OpenStudy (anonymous):

solve 12^x-2= 20 (use log)

OpenStudy (solomonzelman):

"x-2" is the exponent ?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf 12^{x-2}= 20 }\) \(\LARGE\color{blue}{ \bf log~12^{x-2}= log~20 }\) \(\LARGE\color{blue}{ \bf (x-2) log~12= log~20 }\) \(\LARGE\color{blue}{ \bf (x-2) = log~20/log~12 }\) \(\LARGE\color{blue}{ \bf x = 2+log~20/log~12 }\) \(\LARGE\color{blue}{ \bf x = 2+log_{12}~20 }\) can you do it from here ?

OpenStudy (anonymous):

no lol

OpenStudy (anonymous):

i tried

OpenStudy (solomonzelman):

you can use the calculator to solve \(\LARGE\color{blue}{ \bf log_{12}~20 }\) and then finish off by adding the right side, http://www.1728.org/logrithm.htm or do you need exact answer in terms of log ?

OpenStudy (anonymous):

solve for x i need it rounded to the nearest ten thousandth

OpenStudy (anonymous):

thats wrong wolfram